If As demand is p = 10 — X and Bs demand is p = 10 - 2x. The price discriminating monopolists marginal cost is MC(x) = x.
Answer #1

A's demand curve is p=10-x and B's demand curve is p=10-2x

Now, total revenue TR=p*x = 10x-2x2

Then, marginal revenue MR = dTR/dx = 10-4x

Marginal cost for the price discriminating monopolist is MC = x

Then, if the monopolist firm wants B to buy only 2 units of the good, it will maximize its profit at the equilibrium point, MR=MC

or, 10-4x = x

or, 5x=10

or, x=2 units

and p=10-2x = 10-(2*2) = $6 is the equilibrium price for each unit of good x.

Thus, here, only option B is a valid response.